### v2

# A Scalable Bayesian Spatiotemporal Model for Water Level Predictions using a Nearest Neighbor Gaussian Process Approach

Abstract
Obtaining accurate water level predictions are essential for water resource management and implementing flood mitigation strategies. Several data-driven models can be found in the literature. However, there has been limited research with regard to addressing the challenges posed by large spatio-temporally referenced hydrological datasets, in particular, the challenges of maintaining predictive performance and uncertainty quantification. Gaussian Processes (GPs) are commonly used to capture complex space-time interactions. However, GPs are computationally expensive and suffer from poor scaling as the number of locations increases due to required covariance matrix inversions. To overcome the computational bottleneck, the Nearest Neighbor Gaussian Process (NNGP) introduces a sparse precision matrix providing scalability without compromising its inferential capabilities. In this work, we introduce an innovative model in the hydrology field, specifically designed to handle large datasets consisting of a large number of spatial points across multiple hydrological basins, with daily observations over an extended period. We investigate the application of a Bayesian spatiotemporal NNGP model to a rich dataset of daily water level of rivers located in Ireland. The dataset comprises a network of 301 monitoring stations situated in various basins across Ireland, measured over a period of 90 days. The proposed approach allows for prediction of water level at future time points, as well as the prediction at unobserved locations through spatial interpolation. Furthermore, the Bayesian approach provides the benefits in terms of uncertainty propagation and quantification, which are of considerable importance to the hydrology field. Our findings demonstrate that the proposed model outperforms competing scalable approaches in terms of accuracy and precision. Moreover, the proposed model enables predictions across broader range of water bodies with minimal trade-off in predictive performance.
1 Introduction
In Ireland, rivers serve as a primary source of freshwater, providing drinking water to households, supporting industrial processes, and sustaining agriculture. The monitoring of water level enables authorities to predict and manage water availability, especially during dry periods, as low levels can lead to water shortages, while high levels may require preventative measures to avert flooding and contamination in terms of sewage. Accurate predictions of water level are essential for effective water resource management and flood mitigation strategies.
Ireland’s climate contributes significantly to water level fluctuations with heavy precipitation during autumn and winter being a major contributing factor. At these times there is heightened flood risk in many areas. Flood events lead to both direct economic losses, such as damage to infrastructure and property, and indirect losses through disruptions to economic activity. Between 1996 and 2015, the global economic losses from flooding were estimated at $597$ billion USD (Willner et al.,, 2018). Accurate water level prediction is therefore increasingly important in Ireland due to the rising frequency of flood events, as well as to the increasing demand for water across industrial and domestic sectors (Charlton et al.,, 2006). Additionally, the effects of climate change have intensified precipitation patterns, therefore increasing the risks of flooding (Murphy et al.,, 2023). This highlights the critical need for models to accurately predict water level in order to protect communities and support resilient infrastructure planning.
In the hydrology field, water level recordings have been traditionally collected through physical sensors strategically placed along the river stream, creating a network of monitoring stations that collect data over time. The main challenge of making water level predictions is capturing the intrinsic spatiotemporal dynamics of river systems while accounting for related influencing factors. For instance, precipitation events increase water levels while evaporation has the opposite effect, contributing to water level reductions. Indeed, empirical studies have demonstrated that the precipitation in the preceding days are the most significant factor in water level increases (Kenda et al.,, 2020; Ahmed et al.,, 2022).
Another challenge inherent to the task of predicting water level is the quality of the data. Physical sensors are prone to malfunctions, which can result in missing or unreliable measurements that introduce biases into the data. This introduces additional complexity to the prediction task, since data-driven models often require effective pre-processing to handle such issues prior to being able to generate predictions. If not properly managed, poor data quality can significantly compromise the model’s predictive capabilities.
In the literature, two main approaches are employed for predicting water level: hydrological simulations and/or data-driven models. The former uses physical models to simulate the water level dynamics through several obligatory inputs, such as soil and land-use related parameters, leaf area index and snow melt and frost parameters. However, these methods are often computationally costly, especially for large simulations and rely on assumptions of the input parameters derived from published literature (Knijff et al.,, 2010). For the latter, many data-driven methods have been applied to predict water level within a space and time context using machine learning and statistical techniques. Pan et al., (2020) applied a convolutional neural network to predict daily water level over an extended period (30 years) but only considered four water stations. More recently, Pagendam et al., (2023) tackled a large dataset consisting of weekly observations of groundwater depth using a deep neural network with two sub-architectures highlighting the advances of computational efficient methods. However, the mentioned machine learning approaches are limited to predictions only at monitoring sites and often lack the interpretability and transparency as opposed to statistical models. Furthermore, these methods typically do not provide rigorous quantification of predictive uncertainty.
Space-time kriging has been widely applied to predict water level under the frequentest (non-Bayesian) framework (Varouchakis et al.,, 2022; Ruybal et al.,, 2019; Hoogland et al.,, 2010; Júnez-Ferreira and Herrera,, 2013; Kazemi et al.,, 2021). However, the approach suffer from poor scaling as the number of locations increases. The Gaussian Process (GP) model is frequently used in the hydrological literature since it is capable of modelling complex space-time dependent covariance structures while providing a flexible approach for approximating smooth functions. Additionally, the approach has been widely studied under the frequentist and Bayesian inference. There is a strong connection between GPs and regression kriging, and these can be viewed as interchangeable in some scenarios. However, Varouchakis and Hristopulos, (2019) showed that GPs demonstrate superior performance when compared to kriging. In general, both GP and kriging approaches are susceptible to computational limitations when applied to large dataset due to the inversions of covariance matrices.
The seminal paper of Vecchia, (1988) introduced an approximation for the GP that alleviates its computational limitations, making it alternative and feasible approach for large datasets relying on GPs. Based on this concept, Datta et al., 2016a proposed the Nearest Neighbor Gaussian Process (NNGP) model, showing that this approximation is a valid spatial process over the whole spatial domain without compromising its inferential capabilities. More recently, the NNGP approach has been recognised as the state of the art in terms of prediction for spatial and spatiotemporal problems (Abdulah et al.,, 2022). The application of NNGP approach to the field of climate science is a topic already addressed in the existing literature. For instance, (Datta et al., 2016b, ) investigated its predictive power for a particulate matter air quality dataset. Mastrantonio et al., (2019) proposed a hierarchical multivariate NNGP model for predicting monthly precipitation, minimum and maximum temperatures. However, there is little to no consideration in the literature of its application to hydrological datasets.
In the field of hydrology, where missing values are a common occurrence, non-Bayesian approaches, such as those discussed above, have the disadvantage of requiring the use of an auxiliary model to address the missingness of data. In contrast, Bayesian approaches offer a natural solution for imputation through the specification of prior distributions. Molinos et al., (2015) proposed a Bayesian harmonic regression model to predict monthly water level in Irish lakes considering an extended period (1974 - 2012) while accounting for seasonal effects. While this study provided insights into long-term water level trends, it analysed recordings from only 28 spatial locations. Related to water level prediction, Santos-Fernandez et al., (2022) proposed a Bayesian space-time model to predict daily water temperature in rivers considering covariance functions depending on stream and Euclidean distance. Despite the innovation introduced by the authors, computational issues arise again as the method rely on GPs. Moreover, the approach requires prior informations of the dendritic stream network obtained through an extensive processing of geospatial hydrological databases.
Despite the development of spatiotemporal models in the field of hydrology, a few successful studies have successfully addressed the challenge of predicting water level across a dense network of monitoring stations with extensive temporal records. Such studies are even more scarce when considering results encompassing both high predictive performance and computational efficiency. This task remains challenging due to several factors. As it requires an approach capable of simultaneously overcoming the computational limitations, capturing the complex space-time dependencies, accounting for prediction uncertainty and handling data gaps. In order to address this challenge, we propose a Bayesian spatiotemporal NNGP model to predict daily water level considering a large number of stations located in Ireland recorded over an extended period. The dataset consists of a dense network of 301 stations measured over 90 days. The model describe the water level variability in terms of a single predictor (five-days moving average of precipitation) and an additive spatiotemporal effect following a first-order autoregressive process to capture the space-time process. The proposed approach allows for predicting both monitored stations in a future day and predicting unseen locations through spatial interpolation.
The remainder of the paper is structured as follows. In Section 2 we describe the Irish water level and precipitation dataset. Section 3 presents the definition of NNGP, followed by the formulation and detailed explanation of our proposed spatiotemporal NNGP model. In Section 4 we outline the evaluation strategy and performance metrics. Model results are then presented for a synthetic dataset, followed by the application to the Irish water level data. In Section 5 we compare our proposed model with alternative approaches documented in the literature. In Section 6 the limitations of this study are discussed. Finally, Section 7 concludes the paper with a summary of findings and suggestions for future research.
2 Water level and precipitation records in Ireland
The Office of Public Works (OPW) collects water level in daily and sub-daily frequency for an extensive hydrometric network of 301 monitoring stations situated on rivers, lakes and tidal locations throughout the Republic of Ireland. The monitoring stations are primarily located in areas susceptible to flooding or considered critical for water resource management. For this study, we selected only those stations located on rivers and extracted the daily average water level recordings (in meters) for the period from 03/01/2022 to 02/04/2022. The water level measurements correspond to the daily average water depth above the zero level on the local staff gauge. The data was extracted via the Application Programming Interface (API) available at https://waterlevel.ie/ (OPW,, 2021). Given the substantial variability in the water level across rivers and to facilitate the model fitting, we centred the water level measurements by subtracting the mean value at each station. This transformed variable represents the water level relative to the mean stage, which we adopt as the response variable in the analysis. For simplicity, we refer to this variable as water level whenever the context is clear. The water level distribution across the monitoring stations for a sample of days is illustrated in Figure 1. The distances between stations were calculated using the great circle distance to account for the curvature of the earth, thus enhancing precision and ensuring a representation in kilometres.
Figure 1: Water level relative to the mean stage (meters) observed at stations across Ireland between 4 February 2022 and 7 February 2022.
The Irish Meteorological Service, Met Éireann provides forecasting and access to historical data on several weather variables, including 24 hours precipitation accumulation, temperature, wind speed, soil moisture and potential evapotranspiration. However, for this study, only precipitation data were public available for the relevant time frame. We extracted the precipitation variable for the available weather stations, covering the same period as the water level data. The data retrieval was facilitated using the API available at https://marie.ie/repo/met-eireann-api/ (Marie,, 2023).
The observed precipitation did not align spatially with the coordinates of river monitoring stations (see Figure 2). To address this spatial misalignment, the Inverse Distance Weighting (IDW) interpolation was applied to estimate the precipitation at the location of each river station. The IDW is a deterministic interpolation method which the estimated values is computed based on the weighted averages of observed values, where the weights are inversely proportional to the distance between the spatial points. This methodology has been demonstrated to be effective for interpolating precipitation data, as evidenced by Chen and Liu, (2012).
Figure 2: Map of Ireland illustrating the spatial misalignment between river (red) and hydroclimate stations (green).
In order to better capture the relationship between precipitation and water level, particularly with regard to the delayed hydrological response, a five-day moving average was applied to the daily precipitation data. This smoothing technique strengthens the linear correlation with water level fluctuations and helps to capture the rainfall peaks. The resulting smoothed precipitation series was then used as a covariate in the analysis (see Figure 3). For clarity, we refer to this smoothed variable as precipitation throughout the remainder of the manuscript.
Figure 3: Left: relationship between daily precipitation and water level relative to the mean stage. Right: strengthen relationship between precipitation trends based on five-day moving average and water level relative to the mean stage.
As demonstrated in Figure 4, the temporal evolution of water level across the monitoring stations during the study period is illustrated. A common pattern is evident across most stations, with pronounced peaks occurring in early and late February to early March. However, some stations exhibit deviations from this general trend, either with attenuated or delayed peaks.
Figure 4: Temporal evolution of water level for the monitoring stations.
A monthly analysis of water level and precipitation reveals elevated water level in February, primarily attributed to prolonged periods of heavy rainfall. In a similar way, heightened water level are evident in March, aligning with notable precipitation peaks as indicated by the substantial number of outliers observed in the precipitation data (see Figure 5). These precipitation patterns are consistent with the climatology of Ireland, where the wet season typically occurs during the winter months and gradually diminishes as it transitions into spring, contributing to a subsequent decline in water levels.
Figure 5: Distribution of precipitation and water level relative to the mean stage by month.
The impact of water level drivers beyond precipitation is clearly demonstrated when examining the fluctuations in water level across stations over time (Figure 6). Although all six stations experienced peaks in precipitation during late February and early March, the water level response to these events varied notably across stations. This pattern indicates that additional hydroclimate factors such as soil saturation and temperature may be affecting the water level dynamics alongside precipitation.
Figure 6: Water level relative to the mean stage (meters) and precipitation (millimetres) at six river monitoring stations.
3 Methodology
The proposed model draws inspiration by a study on Australian river water temperatures conducted by Santos-Fernandez et al., (2022). This work was selected for its suitability in addressing the challenges posed by long-term spatiotemporal data - specifically, 42 spatial locations measured over 87 days. Notably, the time range closely resembles that of the present work. The authors proposed a Bayesian spatiotemporal model for predicting water temperature which the temporal autocorrelation is captured through an autoregressive parameter and the spatial dependence along the river network is modelled through a Gaussian Process (GP) depending on Euclidean and stream distance. In contrast, our study deals with a significantly larger dataset comprising 301 spatial locations compared to their 42. This substantial increase in spatial points makes the use of a conventional GP computationally impractical due to the inversion of a large spatial covariance matrix. In order to address the computational limitation, we employed the Nearest Neighbor Gaussian Process (NNGP), a scalable alternative that effectively approximates the GP while preserving spatial dependencies. First, we formalise the NNGP approach in Section 3.1. Subsequently, we present the proposed spatiotemporal model in Section 3.2.
3.1 Nearest Neighbor Gaussian Process
Let $\boldsymbol{s}=\{\boldsymbol{s}_{1},\dots,\boldsymbol{s}_{n}\}$ be the set of spatial locations in the spatial domain $\mathcal{S}\subset\mathbb{R}^{d}$ and consider a pure spatial process following a zero-centred GP with covariance matrix $\boldsymbol{\Sigma}=\Sigma(\cdot,\cdot)$, denoted as $\boldsymbol{w}\sim GP(\boldsymbol{0},\boldsymbol{\Sigma})$. The process can be expressed as the product of conditional normal distributions
$p(\boldsymbol{w})=p(w(\boldsymbol{s}_{1}))\prod_{i=2}^{n}p(w(\boldsymbol{s}_{i})\mid w(\boldsymbol{s}_{i-1}),\dots,w(\boldsymbol{s}_{1})).$ (1)
In a setting with a substantial number of spatial locations, a certain elements in the conditional set offer minimal contribution, in such a manner that their exclusion results in negligible loss of information. The underlying idea of the NNGP is to substitute each extensive conditional set by a subset (smaller set) containing at most $m$-nearest neighbors (Vecchia,, 1988). Formally, let $N(\boldsymbol{s}_{i})$ be the subset of at most $m$-nearest neighbors with respect to location $\boldsymbol{s}_{i}$, that is, $N(\boldsymbol{s}_{i})\subset\{\boldsymbol{s}_{1},\dots,\boldsymbol{s}_{i-1}\}$. We have an approximation for Equation 1, known as the Nearest Neighbor Gaussian Process (NNGP) denoted by $\widetilde{\boldsymbol{w}}\sim NNGP(\boldsymbol{0},\widetilde{\boldsymbol{\Sigma}})$
$p(\boldsymbol{w})\approx p(\widetilde{\boldsymbol{w}})=p(w(\boldsymbol{s}_{1}))\prod_{i=2}^{n}p(w(\boldsymbol{s}_{i})\mid\boldsymbol{w}(N(\boldsymbol{s}_{i}))).$ (2)
As demonstrated by Datta et al., 2016a , the stated approximation is a valid multivariate normal distribution and also is a valid spatial process over $\mathcal{S}$. Hence, the predictions are close to the GP model and the NNGP kriging recovers the true spatial process. Defining the neighbor set $N(\boldsymbol{s}_{i})$ requires imposing an order on locations, such that any given order generates a valid joint distribution. Here, we ordered the locations based on the latitude coordinate, as suggested by Datta et al., 2016a .
From probability theory of the multivariate Normal distribution, each of the conditional densities in Equation 2 is also normally distributed with parameters
$\mathbb{E}[w(\boldsymbol{s}_{i})\mid\boldsymbol{w}(N(\boldsymbol{s}_{i}))]=\boldsymbol{a}_{i}\boldsymbol{w}(N(\boldsymbol{s}_{i}))$ (3)
$\mathbb{V}[w(\boldsymbol{s}_{i})\mid\boldsymbol{w}(N(\boldsymbol{s}_{i}))]=d_{i},$ (4)
where $\boldsymbol{a}_{i}=\Sigma(\boldsymbol{s}_{i},N(\boldsymbol{s}_{i}))\Sigma(N(\boldsymbol{s}_{i}),N(\boldsymbol{s}_{i}))^{-1}$ and $d_{i}=\Sigma(\boldsymbol{s}_{i},\boldsymbol{s}_{i})-\boldsymbol{a}_{i}\Sigma(\boldsymbol{s}_{i},N(\boldsymbol{s}_{i}))^{T}$.
The presented NNGP was formulated as a zero-centred distribution but it is trivial to shift the mean to accommodate a non-zero centred process. Furthermore, the process can be simply extended to a space-time setting, with the incorporation of a separable spatiotemporal covariance structure, as detailed in Section 3.2.
3.2 The spatiotemporal model
Let $\boldsymbol{y}_{t}$ be the vector of observed water level for all locations at time $t=1,\dots,T$. Therefore, the stacked vector $[\boldsymbol{y}_{1},\boldsymbol{y}_{2},\dots,\boldsymbol{y}_{T}]$ represents the multivariate distribution for all locations and time points. We propose a model based on the product of conditional distributions
$[\boldsymbol{y}_{1},\boldsymbol{y}_{2},\dots,\boldsymbol{y}_{T}]=p(\boldsymbol{y}_{1}\mid\boldsymbol{X}_{1},\boldsymbol{\theta})\prod_{t=2}^{T}p(\boldsymbol{y}_{t}\mid\boldsymbol{y}_{t-1},\boldsymbol{X}_{t},\boldsymbol{X}_{t-1},\boldsymbol{\theta}),$ (5)
where $\boldsymbol{X}_{t}$ is the design matrix of predictors at time $t$, $\boldsymbol{\theta}$ is the vector of parameters and $\boldsymbol{y}_{1}$ is the distribution at time $t=1$. For the following time points, the distribution is given by
$p(\boldsymbol{y}_{t}\mid\boldsymbol{y}_{t-1},\boldsymbol{X}_{t},\boldsymbol{X}_{t-1},\boldsymbol{\theta})\sim\mathcal{N}(\boldsymbol{y}_{t};\boldsymbol{\mu}_{t},\widetilde{\boldsymbol{\Sigma}}+\tau^{2}\boldsymbol{I}),$ (6)
where $\widetilde{\boldsymbol{\Sigma}}$ is the exponential covariance matrix from a NNGP and $\tau^{2}\boldsymbol{I}$ is the matrix of errors. The mean parameter of the distribution is
$\boldsymbol{\mu}_{t}=\boldsymbol{X}_{t}\boldsymbol{\beta}+\phi(\boldsymbol{y}_{t-1}-\boldsymbol{X}_{t-1}\boldsymbol{\beta}).$ (7)
The model is second-order stationary following a multivariate normal distribution where the time process is an additive term in the mean parameter while the spatial process is included through the spatial covariance matrix. Indeed, the mean at the current time depends on the predictors and the response at the previous time point scaled by the autoregressive parameter $|\phi|<1$. The model presented in Equation 6 exploits the benefits of the normal likelihood by integrating out the spatial process, resulting in what is known as the marginal model (Finley et al.,, 2019). Sampling from this formulation is preferable as it enhances stability by reducing the dimensionality of the parameter space. However, for prediction extrapolating the time dimension, the marginal model is not practical, as it requires the values of predictors for all locations in a future time point. An equivalent formulation, referred to as the latent model, can be used in such cases. This latent model is explicitly expressed in terms of the spatial latent field, that is
$\boldsymbol{y}_{t}=\boldsymbol{X}_{t}\boldsymbol{\beta}+\phi(\boldsymbol{y}_{t-1}-\boldsymbol{X}_{t-1}\boldsymbol{\beta})+\widetilde{\boldsymbol{w}}_{t}+\boldsymbol{\varepsilon}_{t},$ (8)
where $\widetilde{\boldsymbol{w}}_{t}\sim NNGP(\boldsymbol{0},\boldsymbol{\widetilde{\Sigma}})$ and $\boldsymbol{\varepsilon}_{t}\sim\mathcal{N}(\boldsymbol{0},\tau^{2}\boldsymbol{I})$. As stated previously, the marginal model is more computationally efficient, therefore, we used that formulation for sampling from the posterior and for predictions purely in space. In contrast, the latent model (Equation 8) is used for making predictions that require extrapolation over time. Detailed explanations of the predictions are provided in Section 3.3.
Uninformative priors were assumed for the regression parameters, $\mathcal{N}(0,100)$, whereas for the autoregressive parameter a non-informative prior was assigned, $\phi\sim U(-1,1)$. For the variance parameters, we assumed conjugate priors, $\sigma^{2}\sim IG(2,1)$, and a tighter prior for the error variance to ensure the identifiability of the model, $\tau^{2}\sim IG(2,0.1)$. The length-scale parameter was assumed as $l\sim U(0.1,300)$ that corresponds to an effective range of $[0.3,900]$ km. Finally, the number of nearest neighbors was set to 10, as this value produced satisfactory results when tested on a synthetic dataset used to validate the approach, see the results in Section 4.1). This choice aligns with findings by Datta et al., 2016a , who demonstrated that modest values of $m<20$ offers balance between computational efficiency and predictive accuracy in NNGP models.
A full Bayesian inference approach was employed to naturally handle missing values and provide thorough quantification of uncertainty. The model was implemented in Stan using Markov Chain Monte Carlo (MCMC) sampling via No-U-Turn Sampler (NUTS). In order to facilitate the interface with Stan, the cmdstanr package in R was used (Gabry et al.,, 2024). The implementation efficiently completed 500 iterations across 4 chains in approximately 36 minutes. The code used for model implementation is available in this Github repository. All scripts and instructions necessary to reproduce the results are provided.
3.3 Spatiotemporal prediction
The proposed spatiotemporal model allows predictions across both time and space, with the extrapolations being being made along each of the dimensions either individually or simultaneously. First, we present the approach for predicting over space and time at an unobserved location. Following that, we present the method for predicting future values at locations with historical observations. Here, we introduce the notation for the predictive scenarios. Let $[\boldsymbol{y}_{t}]=[\boldsymbol{y_{1}},\dots,\boldsymbol{y}_{t}]$ be the vector of water level measurements for all locations until the observed time $t=1,\dots,T$. Consider an observed location $\boldsymbol{s}_{obs}$ where we want to make predictions at a future time point. Additionally, an unobserved (new) location $\boldsymbol{s}_{new}$ where we want to predict the water level either for a past of future time point. The three prediction scenarios that we implemented are detailed below.
3.3.1 Unobserved (new) location at a past time point
Given an unobserved (new) location denoted by $\boldsymbol{s}_{new}$, we want to obtain the conditional distribution $p(y_{t}(\boldsymbol{s}_{new})\mid[\boldsymbol{y}_{t}])$, for any past time point $t=1,\dots,T$. This can be derived from the joint distribution of the unobserved location and the observed data at the given time point. This derivation is based on the marginal model (Equation 6) by conditioning on the observed data to obtain predictions for unobserved locations. The joint distribution is given by
$\begin{bmatrix}y_{t}(\boldsymbol{s}_{new})\mid y_{t-1}(\boldsymbol{s}_{new})\\ \boldsymbol{y}_{t}(N(\boldsymbol{s}_{new}))\mid\boldsymbol{y}_{t-1}(N(\boldsymbol{s}_{new}))\end{bmatrix}\sim$ $\mathcal{N}$ $\left(\begin{bmatrix}\mu_{t}(\boldsymbol{s}_{new})\\ \boldsymbol{\mu}_{t}(N(\boldsymbol{s}_{new}))\end{bmatrix}\right.,$
$\left.\begin{bmatrix}\sigma^{2}&\Sigma(\boldsymbol{s}_{new},N(\boldsymbol{s}_{new}))^{T}\\ \Sigma(\boldsymbol{s}_{new},N(\boldsymbol{s}_{new}))&\boldsymbol{\Sigma}_{N}\end{bmatrix}\right),$
where $\boldsymbol{\Sigma}_{N}=\Sigma(N(\boldsymbol{s}_{new}),N(\boldsymbol{s}_{new}))+\tau^{2}\boldsymbol{I}$. Notice that the distribution can only be determined by using the preceding distribution of $y_{t}(\boldsymbol{s}_{new})$ until the desired time point time $t$. From probability theory of the multivariate Normal distribution the conditional distribution is
$y_{t}(\boldsymbol{s}_{new})\mid y_{t-1}(\boldsymbol{s}_{new}),\boldsymbol{y}_{t},\boldsymbol{y}_{t-1}\sim\mathcal{N}(\mu_{new},\sigma^{2}_{new})$
$\mu_{new}=\mu_{t}(\boldsymbol{s}_{new})+\Sigma(\boldsymbol{s}_{new},N(\boldsymbol{s}_{new}))(\sigma^{2})^{-1}(\boldsymbol{y}_{t}(N(\boldsymbol{s}_{new}))-\boldsymbol{\mu}_{t}(N(\boldsymbol{s}_{new})))$
$\sigma^{2}_{new}=\sigma^{2}-\Sigma(\boldsymbol{s}_{new},N(\boldsymbol{s}_{new}))\boldsymbol{\Sigma}_{N}^{-1}\Sigma(\boldsymbol{s}_{new},N(\boldsymbol{s}_{new}))^{T}.$
3.3.2 Unobserved (new) location at a future time point
In order to obtain samples of the conditional distribution of $p(y_{t}(\boldsymbol{s}_{new})\mid[\boldsymbol{y}_{t}])$ for a future time point $t>T$, we have to make use of the latent model. This model is necessary because, at future time points, predictors at all other locations are usually unknown, a requirement for the marginal model. However, for the latent model, only the predictors at the target location are needed, though it requires obtaining samples from the spatial latent field specifically for that location, denoted as $\widetilde{w}_{t}(\boldsymbol{s}_{new})$. We proceed by sampling by composition, first drawing from the predictive distribution of $\widetilde{w}_{t}(\boldsymbol{s}_{new})$ and then using the latent model (Equation 8) to draw samples from $y_{t}(\boldsymbol{s}_{new})$. The first step is done by kriging the spatial latent field at the given time point. The joint distribution is
$\begin{bmatrix}\widetilde{w}_{t}(\boldsymbol{s}_{new})\\ \boldsymbol{\widetilde{w}}_{t}(N(\boldsymbol{s}_{new}))\end{bmatrix}\sim\mathcal{N}\left(\boldsymbol{0},\begin{bmatrix}\sigma^{2}&\Sigma(\boldsymbol{s}_{new},N(\boldsymbol{s}_{new}))^{T}\\ \Sigma(\boldsymbol{s}_{new},N(\boldsymbol{s}_{new}))&\widetilde{\boldsymbol{\Sigma}}_{N}\end{bmatrix}\right),$
where $\widetilde{\boldsymbol{\Sigma}}_{N}=\Sigma(N(\boldsymbol{s}_{new}),N(\boldsymbol{s}_{new}))$. Therefore the predictive distribution of $\widetilde{w}_{t}(\boldsymbol{s}_{new})$ has similar form of Equations 3 and 4.
3.3.3 Observed location at a future time point
Given a observed location denoted by $\boldsymbol{s}_{obs}$, we want to obtain the conditional distribution of $p(y_{t}(\boldsymbol{s}_{obs})\mid[\boldsymbol{y}_{t}])$, for a future time point $t>T$. Similar for prediction of unobserved (new) locations at a future time point, this involves sampling from the posterior latent field at the future time point before drawing samples from the distribution of interest. However, the location of interest is known, therefore, kriging is not required here. This is because $\widetilde{w}_{t}(\boldsymbol{s}_{obs})$ is directly part of the spatial latent field $\widetilde{\boldsymbol{w}}_{t}$ eliminating the need for spatial interpolation at the observed location.
4 Application
This section presents an analysis of a synthetic dataset using the proposed model, followed by its application to the Irish water level dataset. The objective of the application to the synthetic data is to evaluate the model performance under controlled conditions, and assess its capability for imputing missing values.
In terms of performance, the goal is to make predictions at locations where there is no data available, as well as to forecast future observations at sites with existing data. In order to evaluate the model performance in both scenarios, an out-of-sample (OOS) evaluation was conducted, splitting the data into two subsets: training and validation. Specifically, the validation dataset consisted of two components: (1) a random selection of stations that included all available time points, and (2) the reservation of the last three days of observations for the remaining stations. The illustrative diagram in Figure 7 represents the data splitting strategy for a hypothetical scenario with 20 stations.
Section 4.1 first presents the data generation procedure with missing values and the application of the mentioned splitting setup. Subsequently, it outlines the results obtained by the proposed model. Section 4.2 presents the validation results of the proposed model for the water level dataset, where 20 stations were randomly selected for testing. This left a training set consisting of 281 stations, each measured over a period of 85 days, resulting in a total of 27,090 observations.
Figure 7: Illustrative diagram of training and validation dataset for out-of-sample evaluation for a scenario with 20 monitoring stations. Three stations were randomly selected to contribute exclusively to the validation dataset, containing all their daily observations. Each tile represents a daily observation and is coloured-coded to indicate whether it belongs to training or validation set.
4.1 Synthetic dataset
We simulated a spatiotemporal dataset composed of 100 spatial locations within a unit square domain, each measured for 20 days, in a total of $100\times 20=2,000$ observations. First we simulated the spatial locations in a unit square. Following that, the spatial covariance matrix was computed using the exponential function with parameters $\sigma^{2}=1$ and $l=0.3$. The response variable was then simulated sequentially from $t=1$ onward, using Equation 6. The mean function (Equation 7) assumed regression coefficients $\boldsymbol{\beta}=[1,3]$ and autoregressive parameter $\phi=0.9$. The parameter values were selected in a manner that ensured the temporal correlation resulted in a long-lasting effect while simultaneously preventing the spatial parameters from being overly smoothed or localised. Finally, the space-time process was fully specified adding the error variance ($\tau^{2}=0.1$) into the spatial covariance. We can summarise the generation of the synthetic dataset as
$\begin{cases}\boldsymbol{x}_{t}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I}),\quad t=1,\dots,20\\ \boldsymbol{s}=\{\boldsymbol{s}_{1},\dots,\boldsymbol{s}_{100}\}\text{ where }\boldsymbol{s}_{i}=(u_{1},u_{2})\text{ and }u_{1},u_{2}\sim Unif(0,1),\quad i=1,\dots,100\\ \boldsymbol{\Sigma}=\Sigma(\boldsymbol{s}_{i},\boldsymbol{s}_{j})=\exp\left(-\frac{|\boldsymbol{s}_{i}-\boldsymbol{s}_{j}|}{0.3}\right)+0.1^{2}\boldsymbol{I}\\ \boldsymbol{y}_{1}\sim\mathcal{N}(1+3\boldsymbol{x}_{1},\boldsymbol{\Sigma}/(1-\phi^{2}))\\ \boldsymbol{\mu}_{t}=1+3\boldsymbol{x}_{t}+0.9\times(\boldsymbol{y}_{t-1}-(1+3\boldsymbol{x}_{t})),\quad t>1\\ \boldsymbol{y}_{t}\sim\mathcal{N}(\boldsymbol{\mu}_{t},\boldsymbol{\Sigma}),\quad t>1.\end{cases}$
In order to assess the imputation capability of the model, 10% of the simulated observations were set as missing. While, for prediction performance, we follow the same setup as presented in Figure 7 but using 10 spatial locations as unobserved locations and leaving out the last day of the remaining observations for validation. The priors for the model were chosen as described in Section 3.2, with an adjustment made for the length-scale parameter to accommodate the spatial distances within the context of this synthetic data. In order to ensure an appropriate prior distribution in a unit square domain, the length-scale parameter was specified as $l\sim U(0.1,1)$. Additionally, the number of nearest neighbors for the NNGP prior was set to 10 to balance computational efficiency and spatial accuracy in the spatial latent field estimation.
Table 1 shows the posterior estimates for the synthetic data. The model successfully produces parameter estimates close to the true parameter values, with the 95% credible interval (CI) capturing all the parameter values, with the exception of the variance parameter. This indicates strong overall model performance in parameter estimation, though minor limitations in estimating the variance of the latent spatial field.
The model demonstrates accurate predictions in both scenarios: (1) predicting the values at previously observed locations for a future day, and (2) estimating the values for unobserved locations on days when other monitoring stations have recorded data. As illustrated in Figure 8, the posterior mean closely aligns the true values with the 95% prediction intervals successfully encompassing 99% of the true values. In terms of imputation capabilities, the model demonstrated high accuracy in recovering the true missing values. As shown in Figure 9, the posterior mean aligns with the true values with a high degree of precision, in fact, the 95% prediction intervals successfully contain the true values for approximately 93% of the missing observations. These results indicate that the proposed model effectively addresses complex spatiotemporal prediction problems where predictions can be across both space and time. Furthermore, the model demonstrates robust performance in imputing missing observations based on the posterior predictive distribution, eliminating the need for an auxiliary imputation model.
Table 1: Posterior estimates of model parameters in the synthetic dataset and its true values.
Parameter True value Mean Median 95% CI
$\beta_{0}$ 1.00 0.58 0.58 (-0.11, 1.22)
$\beta_{1}$ 3.00 2.99 2.99 (2.97, 3.01)
$\phi$ 0.90 0.91 0.91 (0.89, 0.93)
$\sigma^{2}$ 1.00 1.23 1.23 (1.08, 1.41)
$\tau^{2}$ 0.10 0.07 0.07 (0.05, 0.10)
$l$ 0.30 0.25 0.25 (0.21, 0.30)
Figure 8: Predictive performance of the proposed model for the synthetic dataset considering prediction one-day ahead under two different scenarios (unseen and observed locations). The dots represents the posterior mean while the vertical bars represent the 95% prediction interval.
Figure 9: Imputation performance of the proposed model for the synthetic dataset. The dots represents the posterior mean while the vertical bars represent the 95% prediction interval.
4.2 Water level dataset
Table 2 shows the posterior estimates for the water level dataset. The large value of the autoregressive parameter ($\phi=0.87$) highlights the strong temporal dependence between the water level in consecutive days. In other words, the water level in the future days is similar to the previous day if there is no large changes in precipitation. The posterior mean of the length-scale parameter is $l=34.14$ kilometers, which is a considerable range for the spatial decay. This range indicates that locations within approximately 35 kilometers of each other are correlated, which is consistent with the physical continuity of river networks and the gradual change in water levels over short distances. Furthermore, the proportion of variance that is captured by the spatial latent field - $\sigma^{2}/(\sigma^{2}+\tau^{2})\approx 0.76$ emphasising the importance of including the spatial component.
Table 2: Posterior estimates of model parameters in the water level dataset.
Parameter Mean Median 95% CI
$\beta_{0}$0.02 -0.02 (-0.06, 0.02)
$\beta_{1}$ 0.024 0.024 (0.021, 0.027)
$\phi$ 0.878 0.878 (0.874, 0.884)
$\sigma^{2}$ 0.014 0.014 (0.013, 0.014)
$\tau^{2}$ 0.0043 0.0043 (0.0041, 0.0044)
$l$ 34.14 34.14 (32.14, 36.35)
After obtaining samples from the posterior distribution, we used them to derive the posterior predictive distribution of water levels at locations without available observation and at future time points. Figure 10 presents the comparison between the predicted and true water level for forecasts ranging from one to five days ahead. The results demonstrate that the model performs well for short-term predictions, particularly for one and two-day in advance, as the posterior mean is close to the true values with narrow predictive intervals. However, the predictive performance decreases with longer forecast horizons, with the three-day-ahead prediction exhibiting noticeable overestimation and great uncertainty, as reflected by the wider prediction intervals. This increase in uncertainty is a expected result, as longer-term forecasts inherently carry more variability. Similarly, the prediction for an observed location on a future day exhibits superior accuracy and more narrow prediction intervals in comparison to sites without data (see Table 3).
Figure 10: Predictive performance of the proposed model for the water level dataset considering predictions 1 to 5-days ahead under two different scenarios (unseen and observed locations). The dots represents the posterior mean and while the vertical bars represent the 95% prediction interval.
Table 3: RMSE of OOS evaluation for prediction in 1 to 5-days ahead according to data availability (observed and unseen).
Date Observed location Unseen location
29/03/2022 (1-day ahead) 0.035 0.135
30/03/2022 (2-day ahead) 0.073 0.148
30/03/2022 (3-day ahead) 0.101 0.159
01/04/2022 (4-day ahead) 0.131 0.174
02/04/2022 (5-day ahead) 0.153 0.188
Figure 11 depicts the predicted water level - after back-transformation to the original scale - for 20 unobserved locations. At each observed time point, an interpolation was performed to estimate water level at these unseen spatial points. For dates extending beyond the training dataset (29/03/2022 to 02/04/2022), the spatial latent field was first computed for the observed locations and subsequently interpolated in order to predict the water level at the unobserved locations, as outlined in Section 3.3. The limited information beyond 02/04/2022 is reflected in wider prediction intervals, indicating greater uncertainty in predictions.
Overall, the model demonstrated capacity to predict water level at unobserved locations, including for future days. It successfully captured the general temporal patterns of water level variation across most unseen locations. However, for some unobserved locations (stations 18002, 18019, 23030, 25001 and 29002), the predictions were conservative and failed to capture the scale of the variability of water levels, particularly, struggling to accurately predict the peak events. Additionally, at some sites (station 14121 and 22009), the model overestimate the water level fluctuations. These inconsistencies suggest that while the model generally performs well for unobserved locations, further refinement may be necessary to accommodate localised hydrological behaviour and site-specific characteristics.
Figure 11: Predicted water level - after back-transformation to the original scale - and its true values for 20 unobserved stations. The black dots represent the real values, while the blue line represents the posterior mean with the shaded area indicating the 95% prediction interval.
5 Model Comparison
Despite the model’s novelty in the hydrology field, we aim to compare its prediction performance and uncertainty quantification with other methodologies in the literature within two different setups: (1) a comparison with a competitor model for large spatiotemporal datasets and (2) a comparison with models focused on a single basin, a strategy commonly adopted by hydrologists. This approach treats each river system independently, without leveraging information from other basins. For this setup, we also assess the appropriateness of the specification of the proposed model by exploring an alternative formulation for the space-time effect.
The selected criteria for assessing the predictive performance were Root Mean Square Error (RMSE) and Mean Absolute Error (MAE) on the out-of-sample data, which provide a measure of the accuracy of point predictions. In order to evaluate the uncertainty estimation and goodness of fit, the the prediction coverage at the 90% and 95% confidence levels was assessed. This metric measures the proportion of posterior intervals containing the true value, with a good model achieving coverage close to the confidence level. Additionally, the Continuous Ranked Probability Score (CRPS) was calculated, which penalises predictions when the prediction interval fails to capture the true value. Lower CRPS values indicate better performance (Gneiting and Raftery,, 2007).
5.1 Comparison with a competitor model for large spatiotemporal data
For the first comparison, we selected the Gaussian Predictive Process (GPP) model. The model achieves scalability by defining a GP over a smaller set of locations, called knots, and then using kriging to predict the spatial effect at the observed locations (Sahu and Bakar,, 2012). Formally, for a set of knot locations $\boldsymbol{s}^{\star}=\left\{\boldsymbol{s}^{\star}_{1},\dots,\boldsymbol{s}^{\star}_{m}\right\}$ where $m\ll n$, the associated spatial process at time $t$ is denoted by $\boldsymbol{z}_{t}=\left\{z_{t}(\boldsymbol{s}^{\star}_{1}),\dots,z_{t}(\boldsymbol{s}^{\star}_{m})\right\}$. Therefore, the distribution of observations at time $t$ is given by
$p(\boldsymbol{y}_{t}\mid\boldsymbol{X}_{t},\boldsymbol{z}_{t},\boldsymbol{\theta})\sim\mathcal{N}(\boldsymbol{y}_{t};\boldsymbol{X}_{t}\boldsymbol{\beta}+\boldsymbol{\widetilde{w}}_{t},\tau^{2}\boldsymbol{I}),$ (9)
where the spatial effect at observed locations can be approximated by $\widetilde{\boldsymbol{w}}_{t}=\boldsymbol{A}\boldsymbol{\Sigma}_{z}^{-1}\boldsymbol{z}_{t}$. Also, $\boldsymbol{A}$ denotes the covariance between the spatial effect at the observed locations and the knot location and $\boldsymbol{\Sigma}_{z}$ is the covariance matrix of the spatial process $\boldsymbol{z}_{t}$. Moreover, to capture the temporal dependency, we assume the autoregressive model
$\boldsymbol{z}_{t}=\rho\boldsymbol{z}_{t-1}+\boldsymbol{\xi}_{t},\quad\boldsymbol{\xi}_{t}\sim GP(\boldsymbol{0},\boldsymbol{\Sigma_{z}})$ (10)
The set of knots were placed using the function makegrid from the R package sp (Pebesma and Bivand,, 2005). A regular grid with approximately 100 points was initially created across the study area, and only the points that fell inside the study boundaries were retained, resulting in a total of 104 knots. The implementation of the GPP model used for comparison is available in the R package spTimer (Bakar and Sahu,, 2015).
Table 4: Comparison of performance between the proposed NNGP model and another scalable model for large spatiotemporal data (GPP).
Model Coverage CRPS RMSE MAE
95% 90% space space-time time space space-time time
NNGP 95.1 93.3 1.66 0.22 0.17 0.12 0.15 0.14 0.09
GPP 95.8 93.9 2.71 0.31 0.27 0.25 0.22 0.23 0.21
The results of model comparison between the proposed model (NNGP) and the competitor scalable spatiotemporal model (GPP) are presented in Table 4. The proposed model demonstrated lower error in terms of RMSE and MAE for the three predictive scenarios, including unobserved locations at observed times, monitoring stations at future times and unobserved locations at future times. Specifically, the NNGP model has greater predictive accuracy in both spatial and temporal extrapolations. Additionally, the NNGP exhibited lower CRPS values, which reflect narrower and more reliable prediction intervals and achieved better coverage closer to ideal values. These findings highlight that the NNGP outperforms the GPP model both in predictive accuracy and uncertainty quantification, making it a more effective approach for addressing large-scale spatiotemporal hydrological problems.
5.2 Comparison with a single-basin model
In this section, we investigate the advantages of the proposed model by comparing it with models that focus on a single basin, as well adopted in the hydrology field (Varouchakis et al.,, 2022; Varouchakis and Hristopulos,, 2019; Kazemi et al.,, 2021; Ruybal et al.,, 2019). Single-basin models are convenient because they alleviate the computational bottleneck of inverting large covariance matrices while preserving the unique characteristics of the basin. However, they may limit the model’s ability to leverage information from multiple basins, which could enhance its performance. Our comparison aims to highlight the trade-off between computational feasibility and model robustness when scaling to larger, multi-basin datasets.
A secondary objective of this comparison is to investigate the benefits of incorporating the autoregressive process into the dependent variable, as proposed in this work (Equations 6 and 7) compared to embedding it within the spatial process (Equations 9 and 10). While both models capture the spatial dependence, they use different approaches - using nearest neighbors to condition each location to small conditional sets for each location (NNGP) and applying dimension reduction followed by kriging (GPP) - which may lead to an unfair comparison when investigating the placement of the autoregressive component. To ensure a more balanced evaluation, we instead model the spatial process assuming a standard GP prior in both formulations. These models are presented in Equations 11 and 12. This comparison allows us to assess how the placement of the autoregressive component affects a model when dealing with spatiotemporal data collected over extended period of time.
In terms of the basin district to be analysed, the Shannon basin was selected for the single-basin models due to its hydrological importance, covering approximately one fifth of Ireland’s land area (see Figure 12). Based on the training and validation split outlined in Section 4, we selected the stations within Shannon, yielding 94 stations for training and 7 for validation. To enable a direct comparison between the the proposed model and single-basin models, the performance was evaluated exclusively at these stations within the Shannon.
Figure 12: River basin districts of Ireland and the Shannon basin boundaries highlighted in black chosen as reference basin for comparison with single-basin models.
GP 1: Autoregressive process into the dependent variable
Spatiotemporal GP model with mean parameter at current time depends on the response in the previous time.
$p(\boldsymbol{y}_{t}\mid\boldsymbol{y}_{t-1},\boldsymbol{X}_{t},\boldsymbol{X}_{t-1},\boldsymbol{\theta})\sim\mathcal{N}(\boldsymbol{y}_{t};\boldsymbol{\mu}_{t},\boldsymbol{\Sigma}+\tau^{2}\boldsymbol{I})$
$\boldsymbol{\mu}_{t}=\boldsymbol{X}_{t}\boldsymbol{\beta}+\phi(\boldsymbol{y}_{t-1}-\boldsymbol{X}_{t-1}\boldsymbol{\beta}).$ (11)
GP 2: Autoregressive process within the spatial process
Spatiotemporal GP model with the temporal dependence is incorporated into the spatial process.
$p(\boldsymbol{y}_{t}\mid\boldsymbol{X}_{t},\boldsymbol{z}_{t},\boldsymbol{\theta})\sim\mathcal{N}(\boldsymbol{y}_{t};\boldsymbol{X}_{t}\boldsymbol{\beta}+\boldsymbol{w}_{t},\tau^{2}\boldsymbol{I})$
$\boldsymbol{w}_{t}=\rho\boldsymbol{w}_{t-1}+\boldsymbol{\xi}_{t},\quad\boldsymbol{\xi}_{t}\sim GP(\boldsymbol{0},\boldsymbol{\Sigma}).$ (12)
Table 5: Comparison of performance between the proposed NNGP model and single-basin models (GP 1 and GP 2).
Model Coverage CRPS RMSE MAE
95% 90% space space-time time space space-time time
NNGP 95.81 93.97 1.40 0.18 0.16 0.14 0.12 0.14 0.10
GP 1 95.76 94.17 0.99 0.18 0.15 0.12 0.12 0.13 0.08
GP 2 98.52 98.02 1.50 0.26 0.12 0.16 0.20 0.10 0.13
The results of model comparison between the proposed NNGP model, which uses multi-basin data and the single-basin models GP 1 and GP 2 are shown in Table 5. Comparing the two GP models, the GP 1 model exhibits substantially lower RMSE and MAE compared to GP 2 when partial historical information is available for the observation to be predicted - either at unobserved locations with observed times or at monitoring stations projecting into future time points. However, in the absence of such historical data, a scenario that requires extrapolation across both space and time, the GP 2 model performs better in terms of both error metrics. This suggests that incorporating the autoregressive process into the spatial latent field (as in GP 2) provides a more effective mechanism for capturing the latent spatiotemporal dynamics when historical observations are unavailable. These findings highlight the strengths of each model depending on the predictive scenario, with GP 1 better suited for cases where the prediction is purely spatial or temporal. While, GP 2 excelling in capturing latent spatial-temporal dependencies when data is more limited.
In terms of the uncertainty estimation, the GP 1 model showed smaller CRPS compared to GP 2, indicating that the prediction intervals are narrower and more concentrated around the observed values. Additionally, the posterior predictive distribution for GP 2 is more dispersed as shown by the higher coverage values above the nominal level. These results suggest that incorporating the autoregressive process into the dependent variable (as in GP 1 and in the proposed NNGP model) is more appropriate for capturing temporal dependencies in long-term spatiotemporal datasets. Moreover, in that formulation, the autoregressive parameter offers a more natural interpretation, as it is directly tied to the response variable, specifically quantifying the degree of dependence between water levels on successive days.
When compared the performance between the proposed NNGP model and the GP 1 model, which has the same formulation but assumes a GP prior for the spatial latent field, the single-basin GP 1 model exhibited slightly superior predictive accuracy. While the GP 1 model achieved marginally lower RMSE and MAE values, the differences were not substantial. In terms of uncertainty quantification, the 95% coverage for both NNGP and GP 1 models were close to the ideal level, while the 90% coverage for all the models exceeded the target. However, the NNGP model’s coverage was slightly closer to the desired level, indicating it provides more reliable uncertainty estimates. This is particularly important for hydrological applications where the range of the predicted water level is crucial for risk assessment and decision-making.
The results demonstrate that the NNGP approximation exhibits comparable predictive performance to its full GP method. Additionally, the NNGP balances between predictive accuracy and precision. While there is a minimal trade-off in performance compared to the single-basin model, GP 1, the key advantage of the NNGP lies in its scalability and its ability to extend predictions across a broader range of river basins, making it more versatile for large-scale hydrological applications.
6 Limitations
This work proposes a model for predicting water level on rivers in Ireland considering a long-term spatiotemporal data. This approach has been demonstrated to be effective, as evidenced by the promising results. However, several limitations should be noted. As previously mentioned, the model incorporates only precipitation as a predictor. The inclusion of additional hydroclimate predictors such as temperature, wind speed, soil moisture and potential evapotranspiration would potentially improve the predictive performance and reliability of the model. Additionally, the spatial misalignment problem between the observed precipitation, and the water level measurements was addressed using Inverse Distance Weighting interpolation. Despite the simplicity and efficacy of this methodology, more sophisticated probabilistic interpolation could be investigated to potentially enhance the precision of estimations. Nevertheless, the development of such techniques is not the focus of this study, as the misalignment problem introduces additional methodological and computational complexities, and constitutes a research area in its own right (Gotway and Young,, 2002; Finley et al.,, 2014).
In the proposed model, the spatial dependence is captured through a spatial latent field specified as a NNGP for a given covariance function, which fundamentally depends on the Euclidean distance. The introduced structure is straightforward and effective, ensuring computational feasibility without compromising performance. However, this formulation simplifies the complex spatial dependencies inherent in river networks, where flow directionality, varying water volume and confluences may impact the water level measurements along the river. Covariance models that explicitly account for these hydrological features have been developed by Hoef et al., (2006); Cressie et al., (2006); Hoef and Erin, (2010). However, implementing these models is labour-intensive and time-consuming, as it requires extensive processing of stream network data, often involving manual steps. This challenge is further intensified in geographical areas characterised by multiple basins and complex stream connections, such as those found in Ireland.
7 Conclusion
In this paper, we proposed a spatiotemporal NNGP model for predicting water level in a long-term dataset across both space and time. This model not only delivers accurate predictions but also quantifies uncertainty for both monitored stations and unobserved locations. By accounting for spatiotemporal dependencies and incorporating precipitation information, the model offers insights of the impact of precipitation on water levels over time, thereby supporting more informed decision-making in flood mitigation strategies and resource planning.
Although the main attention in water level prediction in Ireland is predicting for the monitoring location at future time point, since they are strategically placed in areas susceptible to flooding or critical for hydrological control, there is considerable interest in extending predictions to surrounding, unmonitored regions.
We compared our approach with the GPP model, a competitor model for large spatiotemporal data well known in the literature. The proposed model demonstrated superior performance in terms of prediction accuracy and a higher degree of confidence its estimations. Furthermore, we also compared our proposed model with approaches encountered in the field of hydrology. Typically, these approaches rely on GP models for a single basin, thus, avoiding the computational limitations of inversions of large covariance matrices. Despite the minimal loss in performance, the NNGP offers the advantage of handling a large number of spatial points across multiple basins. This scalability enables broader application to large-scale river networks, extending prediction capabilities across a wider range of river basins, and capturing the interconnected spatial dependencies within hydrological systems.
Furthermore, our findings demonstrated that for large spatiotemporal datasets with observations collected over an extended period, the specification of the time component in the proposed model enhances the predictive accuracy in scenarios focused exclusively on spatial or temporal predictions. The advantages of the specified temporal component were also observed in the precision of the estimations, with the predictions closer to the true values and narrower prediction intervals. This finding indicates that the given specification of the temporal component is a crucial factor in the successful solution to the given hydrological problem.
Despite the importance of the specification of a temporal parameter, the key component of our model is the inclusion of a NNGP spatiotemporal latent field. The NNGP overcomes the computational challenges of a traditional GP when analysing a large dataset. While the methodology is still under active research, this work introduces a novelty in the application of the NNGP specifically for predicting water levels in river networks. Moreover, existing models in the hydrology literature have not explored a Bayesian statistical model accounting for a large dataset that span both spatial and temporal dimensions.
As a future research direction, the model could be extended by applying the sparsity-inducing NNGP prior not only to the spatial latent field defined over Euclidean space, but also to the covariance structures for stream networks presented in Hoef and Erin, (2010). The proposed extension would improve the modelling of complex spatial autocorrelation along stream segments and allow Bayesian inference for large hydrological datasets without compromising its predictive performance. To the best of our knowledge, this approach has not yet been explored in the literature.
8 Acknowledgments
This publication has emanated from research conducted with the financial support of Science Foundation Ireland under Grant number 18/CRT/6049 and the jointly funded by Taighde Éireann - Research Ireland, and GSI under Grant number 20/FFP-P/8610.